The number of alpha and beta particle emitted in the following radioactive decay 200X90-168Y80
 
 

Grade:12

1 Answers

Arun
4 years ago
On emission of one α-particle, atomic number decreases by 2 units and atomic mass number decrease by 4 units.  Here, decrease in mass number = 200 - 168 = 32 Number of α-particles = 32/4 = 8.  While the emission of β-particle does not effect the mass number and atomic number increases by 1 unit. Number of β-particles = 16 - 10 = 6.